OpenStudy (anonymous):

Write a proportion that could be used to solve for each variable. Then solve 60 pens in 25 boxes 12 pens in y boxes

OpenStudy (anonymous):

pens over boxes equals pens over boxes \[\frac{60}{25}=\frac{12}{y}\] or boxes over pens equal boxes over pens \[\frac{25}{60}=\frac{y}{12}\] or boxes over boxes equals pens over pens \[\frac{y}{25}=\frac{12}{60}\]

OpenStudy (anonymous):

doesn't matter what ratio you use, you get the same answer \[y=\frac{25\times 12}{60}\]